How many gramz and atoms of Oxygen are there in 15 grams of sodium bicarbonate(Molar mass of NaHCO3=84.01g/mol)


1
Expert's answer
2020-08-02T15:49:41-0400

Proportion 1:

84.01 – 100 %

(16×3) – x %

x = 57.13 % - fraction of Oxygen in NaHCO3

Proportion 2:

15 g – 100 %

y – 57.13 %

y = 8.57 g - mass of Oxygen in 15 g of NaHCO3

A mole of oxygen atoms (with the formula O) contains 6.02×1023 O atoms.

In this problem there are (8.57 / 16) = 0.535 moles of Oxygen atoms.

Total number of Oxygen atoms = 0.535×6.02×1023 = 3.22×1023 atoms.

Answer: 8.57 g and 3.22×1023 atoms